OpenStudy (anonymous):

Lim x-> -3 (x^3+27)/[Sqrt(x^2 + 7) - 4]

OpenStudy (anonymous):

just substitute -3 on x

OpenStudy (anonymous):

and if is possible factorize first!

OpenStudy (anonymous):

if there is no simplest form of f(x) is gonna be undefined

OpenStudy (anonymous):

\[\lim_{x \rightarrow -3} \frac{ (x^3+27) }{ \sqrt{x^2 + 7} - 4 }\] ans : -36

OpenStudy (tkhunny):

Why? It looks a little magic at this point. Have you considered either rationalizing the denominator or can we use l'hospital's rule?

OpenStudy (anonymous):

we cannot use L'hospital

OpenStudy (tkhunny):

Okay, since that was the SECOND option, how about trying the FIRST option?

OpenStudy (anonymous):

\[\lim_{x \rightarrow -3} \frac{ (x+3) (x^2-3 x+9) (\sqrt{x^2 + 7} + 4)}{ (\sqrt{x^2 + 7} - 4)(\sqrt{x^2 + 7} + 4) }\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow -3} \frac{ (x+3) (x^2-3 x+9) (\sqrt{x^2 + 7} + 4)}{ (x+3)(x-3) }\]

OpenStudy (anonymous):

@tkhunny is this correct ?

OpenStudy (anonymous):

the hospital rule is the one that is necessary to divide all the eq. by the largest exponext of x?

OpenStudy (anonymous):

nope sorry that come when you have infinity

OpenStudy (tkhunny):

Let's see... The denominator \((\sqrt{x^{2}+7} - 4)(\sqrt{x^{2}+7} + 4) = (x^{2} + 7) - 16 = x^{2} - 9 = (x+3)(x-3)\) Okay. The numerator also looks good. Where does that leave us? Since we are NOT substituting x = -3, it seems \(\dfrac{x+3}{x+3} = 1\) and the limit should be easier after that reduction.

OpenStudy (anonymous):

but i am not getting -36 @tkhunny

OpenStudy (anonymous):

ok got it ... thanks a lot !!

OpenStudy (tkhunny):

After getting rid of the (x+3)'s, the remaining function is continuous at x = -3 and you can just substitute x = -3. Be careful enough and you will get it. Good work.